My questions are. How do you write an expression for.
1.)nince less than twice a number equals the same number increased by four.
2.)six less than five times a number is twenty-nine.
It has been a very long time since I have been in school& trying to figure out how to write expressions is very confusing to me.
Thanks!

Answer by piglet162431(90)

1.2n-9=n+4
2.5n-6=29
Note: don't write expression copying word by word in ORDER. Sometimes for subtracting you might be subtracting, or subtracting FROM.